OpenStudy (anonymous):

complete the square x^2+2x=3

OpenStudy (anonymous):

x = -1 x = 3

OpenStudy (wolf1728):

Get the coefficient of x which is 2, divide it by 2, then square it, then add it to both sides x^2 + 2x + 1 = 3 +1 (x+1)^2 = 4 Then taking the square root of both sides x +1 =2 x = 1 x = -3

OpenStudy (wolf1728):

(the other root "-3" is obtained because the other square root of 4 is -2) x + 1 = -2 x = -3